\begin{thebibliography}{99}




\bibitem{Bekenstein}
J. D. Bekenstein, Transcendence of the law of baryon-number conservation in black hole physics, Phys. Rev. Lett. 28, 452 (1972).




\bibitem{Chase}
J. E. Chase, Event horizons in Static Scalar-Vacuum Space-Times, Commun. Math. Phys. 19, 276 (1970).



\bibitem{C. Teitelboim}
C. Teitelboim, Nonmeasurability of the baryon number of a black-hole, Lett. Nuovo Cimento 3, 326 (1972).





\bibitem{Ruffini-1}
R. Ruffini and J. A. Wheeler, Introducing the black hole, Phys. Today 24, 30 (1971).




\bibitem{Hod-2}
S. Hod,Stationary resonances of rapidly-rotating Kerr black holes, The Euro. Phys. Journal C 73, 2378 (2013).




\bibitem{Hod-5}
S.Hod, The superradiant instability regime of the spinning Kerr black hole, Phys. Lett. B 758, 181(2016).



\bibitem{c1}
Carlos Herdeiro,Vanush Paturyan,Eugen Radu, D.H. Tchrakian,Reissner-Nordstr$\ddot{o}$m black holes with non-Abelian hair,
Phys. Lett. B 772(2017)63-69.

\bibitem{c2}
Maur\'{\i}cio Richartz, Carlos A. R. Herdeiro, Emanuele Berti,
Synchronous frequencies of extremal Kerr black holes: resonances, scattering and stability,Phys. Rev. D 96, 044034 (2017).

\bibitem{c3}
Yves Brihaye, Carlos Herdeiro, Eugen Radu, D.H. Tchrakian,
Skyrmions, Skyrme stars and black holes with Skyrme hair in five spacetime dimension,JHEP 1711(2017)037.




\bibitem{Benone}
C. L. Benone, L. C. B. Crispino, C. Herdeiro, and E. Radu, Kerr-Newman scalar clouds, Phys.
Rev. D 90, 104024 (2014).






\bibitem{Herdeiro-2}
C. Herdeiro, E. Radu, and H. R\'{u}narsson, Non-linear QQ-clouds around Kerr black holes, Phys. Lett. B 739, 302 (2014).






\bibitem{mr3}
Elizabeth Winstanley,Classical Yang-Mills black hole hair in anti-de Sitter space, Lect.Notes Phys.769:49-87,2009.



\bibitem{NSHB}
Yan Peng, Hair mass bound in the black hole with non-zero cosmological constants,
Physical Review D 98,104041,(2018).




\bibitem{NSHB1}
Yan Peng,Hair distributions in noncommutative Einstein-Born-Infeld black holes,Nucl. Phys. B 941(2019)1-10.







\bibitem{Yves Brihaye}
Yves Brihaye, Thomas Delplace, Carlos Herdeiro, Eugen Radu,
An analytic effective model for hairy black holes,
Phys. Lett. B 782(2018)124-130.



\bibitem{Rogatko2}
Marek Rogatko,Uniqueness of higher-dimensional phantom field wormholes,
Phys. Rev. D 97(2018)no.2,024001.





\bibitem{Degollado}
J. C. Degollado and C. A. R. Herdeiro, Stationary scalar configurations around extremal charged black holes,
Gen. Rel. Grav. 45, 2483 (2013).


\bibitem{Cunha}
P. V. P. Cunha, C. A. R. Herdeiro, E. Radu, and H. F. R\'{u}narsson, Shadows of Kerr black holes with scalar hair, Phys.
Rev. Lett. 115, 211102(2015).






\bibitem{Brihaye}
Y. Brihaye, C. Herdeiro, and E. Radu, Inside black holes with synchronized hair, Phys. Lett. B 760, 279 (2016).


\bibitem{Bekenstein-1}
J. D. Bekenstein, Black hole hair: 25-years after, arXiv:gr-qc/9605059.


\bibitem{CAR}
Carlos A. R. Herdeiro, Eugen Radu,Asymptotically flat black holes with scalar hair: a review, Int.J.Mod.Phys.D
24(2015)09,1542014.









\bibitem{Hod-6}
S.Hod, No-scalar-hair theorem for spherically symmetric reflecting stars, Physical Review D 94, 104073 (2016).

\bibitem{Hod-7}
S.Hod, No nonminimally coupled massless scalar hair for spherically symmetric neutral reflecting stars,Physical Review D 96, 024019 (2017).



\bibitem{Bhattacharjee}
Srijit Bhattacharjee, Sudipta Sarkar, No-hair theorems for a static and stationary reflecting star, Physical Review D 95, 084027 (2017).









\bibitem{Hod-8}
S.Hod, Charged massive scalar field configurations supported by a spherically symmetric
charged reflecting shell,Physics Letters B 763, 275 (2016).



\bibitem{Hod-9}
S.Hod, Marginally bound resonances of charged massive scalar fields in the background of
a charged reflecting shell, Physics Letters B 768(2017)97-102.


\bibitem{Yan Peng-1}
Yan Peng, Bin Wang, Yunqi Liu, Scalar field condensation behaviors around reflecting shells in
Anti-de Sitter spacetimes, Eur.Phys.J. C 78 (2018) no.8, 680.


\bibitem{Yan Peng-2}
Yan Peng, Scalar field configurations supported by charged compact reflecting stars in a curved spacetime,
Physics Letters B 780(2018)144-148.



\bibitem{Hod-10}
Shahar Hod, Charged reflecting stars supporting charged massive scalar field configurations, European Physical
Journal C 78, 173 (2017).











\bibitem{Yan Peng-3}
Yan Peng, Static scalar field condensation in regular asymptotically AdS reflecting star backgrounds,
Phys. Lett. B 782(2018)717-722.



	
\bibitem{Yan Peng-4}
Yan Peng, On instabilities of scalar hairy regular compact reflecting stars, JHEP 10(2018)185.


\bibitem{Yan Peng-5}
Yan Peng, Hair formation in the background of noncommutative reflecting stars, Nucl.Phys. B 938(2019)143-153.





\bibitem{cond1}
Filipp N. Rybakov, Egor Babaev, The absence of superconductivity in the next-to-leading order Ginzburg-Landau
functional for Bardeen-Cooper-Schrieffer superconductor, arXiv:2106.0263.





\bibitem{cond2}
A. V. Vagov, A. A. Shanenko, M. V. Milo\v{s}evi\'{c},
V. M. Axt, and F. M. Peeters, Extended ginzburg-landau formalism: Systematic expansion in
small deviation from the critical temperature,
Phys.Rev.B 85,014502(2012).





\bibitem{cond3}
A. Vagov, A. A. Shanenko, M. V. Milo\v{s}evi\'{c}, V. M. Axt, V. M. Vinokur, J. A. Aguiar,
and F. M. Peeters, Superconductivity between standard types: Multiband versus
single-band materials, Phys. Rev. B 93,174503(2016).








\bibitem{dg}
Yun Soo Myung, Yong-Wan Kim, Young-Jai Park,
Thermodynamics and phase transitions in the Born-Infeld-anti-de Sitter black holes,
Phys. Rev. D 78(2008)084002.



\bibitem{sh}
Ahmad Sheykhi, Doa Hashemi Asl, Amin Dehyadegari,
Conductivity of higher dimensional holographic superconductors with nonlinear electrodynamics,
Phys. Lett. B 781(2018)139.






\bibitem{Rogatko1}
Bart{\l}omiej Kiczek, Marek Rogatko, Karol I. Wysokinski,
Holographic DC SQUID in the presence of dark matter, JCAP 01(2021)063.









\bibitem{mr1}
D. N$\acute{u}\tilde{n}$ez, H. Quevedo, and D. Sudarsky,Black Holes Have No Short Hair, Phys. Rev. Lett.
76, 571 (1996).



\bibitem{mr2}
S. Hod,Hairy Black Holes and Null Circular Geodesics,Phys. Rev. D 84, 124030 (2011).


\bibitem{fc}
Yan Peng, Upper bound on the radii of regular ultra-compact star photonspheres,
Phys. Lett. B 790(2019)396-399.



\bibitem{Basu}
Pallab Basu, Pankaj Chaturvedi, Prasant Samantray,
Chaotic dynamics of strings in charged black hole backgrounds, Phys. Rev. D 95(2017)066014.






\bibitem{Rogatko}
Bart{\l}omiej Kiczek, Marek Rogatko,Ultra-compact spherically symmetric dark matter charged star objects,
JCAP 09(2019)049.




\bibitem{Peng Wang}
Peng Wang, Houwen Wu, Haitang Yang,
Thermodynamics and Phase Transition of a Nonlinear Electrodynamics Black Hole in a Cavity,
JHEP 07(2019)002.












\end{thebibliography}

\documentclass[aps,showpacs,preprintnumbers,amsmath, amssymb]{revtex4}

\usepackage{float}
\usepackage{graphics,epsfig}
\usepackage{graphicx}
\usepackage{dcolumn}
\usepackage{bm}

\begin{document}
\baselineskip=0.8 cm

\title{{\bf No scalar condensations outside reflecting stars with coupling terms from Ginzburg-Landau models }}
\author{Guohua Liu$^{1}$\footnote{liuguohua1234@163.com}}
\author{Yan Peng$^{2}$\footnote{yanpengphy@163.com}}
\affiliation{\\$^{1}$ School of Physics and Physical Engineering, Qufu Normal University, Qufu, Shandong, 273165, China}
\affiliation{\\$^{2}$ School of Mathematical Sciences, Qufu Normal University, Qufu, Shandong 273165, China}

\vspace*{0.2cm}
\begin{abstract}
\baselineskip=0.6 cm
\begin{center}
{\bf Abstract}
\end{center}

We consider static scalar fields coupled to the gradient
where the coupling also appears in next-to-leading
order Ginzburg-Landau models.
We study condensation behaviors of scalar fields outside
regular compact reflecting stars in the asymptotically flat background.
For non-negative coupling parameters, we prove that the
reflecting star cannot support coupled static scalar fields.

\end{abstract}

\pacs{11.25.Tq, 04.70.Bw, 74.20.-z}\maketitle
\newpage
\vspace*{0.2cm}

\section{Introduction}

The famous black hole no hair theorem has attracted lots
of interest from mathematicians and physicists.
It states that static scalar fields cannot exist outside
horizons of black holes in the asymptotically flat
background, for progress see references \cite{Bekenstein}-\cite{Brihaye}
and reviews \cite{Bekenstein-1,CAR}. It seems natural since
horizons can absorb scalar fields outside black holes.
However, such no scalar field property
also appears in horizonless spacetime.

In the horizonless asymptotically flat background,
Hod recently found that neutral compact reflecting stars
cannot support formation of scalar field hairs \cite{Hod-6}.
It was further found that such no hair property appears
for asymptotically flat neutral compact reflecting stars and scalar fields
nonminimally coupled to the curvature \cite{Hod-7}.
This no scalar hair theorem was also extended to
neutral compact reflecting stars in the asymptotically dS spacetime \cite{Bhattacharjee}.
It was further found that large charged compact reflecting stars
cannot allow the existence of scalar fields \cite{Hod-8}-\cite{Yan Peng-5}.
On the other side, in the extended Ginzburg-Landau superconductor model, there is
a new term with scalar fields coupled to the gradient \cite{cond1,cond2,cond3}.
So it is of some interest to generalize the discussion in \cite{Hod-6}
by considering scalar fields coupled to the gradient.

This work is organized as follows. We firstly introduce the
horizonless compact reflecting star gravity system with
exterior scalar fields coupled to the gradient.
For a non-negative coupling parameter, we prove that
horizonless compact reflecting star cannot support
the existence of static scalar field hairs. At last,
we present our main conclusions.

\section{No coupled scalar fields in the horizonless reflecting star background}

According to extended Ginzburg-Landau models, the scalar fields
may be coupled to the gradient \cite{cond1,cond2,cond3}.
We are interested in the case of coupled scalar fields
in the asymptotically flat spacetime, whose Lagrangian density
is given by \cite{dg,sh,Rogatko1}
\begin{eqnarray}\label{lagrange-1}
\mathcal{L}=R-|\nabla_{\alpha} \psi|^{2}-\xi\psi^{2}|\nabla_{\alpha} \psi|^{2}-V(\psi^{2}).
\end{eqnarray}
Here we label R as the Ricci curvature and $\psi=\psi(r)$ as the scalar field.
And $\xi$ is the parameter describing coupling strengthen between
scalar fields and the gradient. In order to obtain the final conclusion,
we assume $\xi>0$ in this work. The scalar field potential
$V(\psi^2)$ satisfying $V(0)=0$ and $\dot{V}=\frac{dV(\psi^2)}{d(\psi^2)}> 0$,
which is apparently satisfied by free scalar fields with $V(\psi^2)=\mu^2\psi^2$
and mass $\mu$.

The exterior spacetime of spherically symmetric horizonless compact star
is \cite{mr1,mr2,fc,Basu,Rogatko,Peng Wang}
\begin{eqnarray}\label{AdSBH}
ds^{2}&=&-ge^{-\chi}dt^{2}+\frac{dr^{2}}{g}+r^{2}(d\theta^2+sin^{2}\theta d\phi^{2}).
\end{eqnarray}
The metric functions $\chi$ and $g$ only depend on the
radial coordinate r. The angular coordinates are
labeled as $\theta$ and $\phi$ respectively.
And the radius of the compact star is labeled as $r_{s}$.

With the Lagrangian density (1), we obtain the equation of motion
\begin{eqnarray}\label{BHg}
(1+\xi\psi^2)\psi''+[(1+\xi\psi^2)(\frac{2}{r}-\frac{\chi'}{2}+\frac{g'}{g})+2\xi \psi \psi']\psi'-(\xi \psi'^{2}+\frac{\dot{V}}{g})\psi=0.
\end{eqnarray}

At the star radius, we impose the reflecting condition for the scalar field as
\begin{eqnarray}\label{BHg}
\psi(r_{s})=0.
\end{eqnarray}

At spatial infinity, metric functions have following asymptotical behaviors
\begin{eqnarray}\label{AdSBH}
\chi\rightarrow 0,~~~~~~g\rightarrow 1~~~~~~for~~~~~~r\rightarrow \infty.
\end{eqnarray}

The energy density can be expressed as
\begin{eqnarray}\label{BHg}
\rho=-T^{t}_{t}=g\psi'^{2}+g\xi\psi^{2}\psi'^{2}+V(\psi^{2}).
\end{eqnarray}

Physically acceptable solution requires that gravitational mass
$M=\int_{0}^{\infty}4\pi r^{2}\rho dr$ is finite, from which
we obtain the following relation
\begin{eqnarray}\label{BHg}
r^{2}\rho\rightarrow 0~~~~~for~~~~~r\rightarrow \infty.
\end{eqnarray}

At the infinity, relations (6), (7) and $\xi>0$ yields the condition
\begin{eqnarray}\label{InfBH}
&&\psi(\infty)=0.
\end{eqnarray}

According to relations (4) and (8), the scalar field $\psi(r)$ must possess at least
one extremum point ${r}_{peak}$, where the function reaches a positive local maximum
value or a negative local minimum value \cite{Hod-6}.
In the case that $\psi(r)$ reaches a positive local maximum value,
the scalar field satisfies the following relation
 \begin{eqnarray}\label{InfBH}
\{ \psi>0,~~~~\psi'=0~~~~and~~~~\psi''\leqslant0\}~~~~for~~~~r={r}_{peak}.
\end{eqnarray}

From relation (9) and $\xi>0$, we deduce the characterized inequality
\begin{eqnarray}\label{BHg}
(1+\xi\psi^2)\psi''+[(1+\xi\psi^2)(\frac{2}{r}-\frac{\chi'}{2}+\frac{g'}{g})+2\xi \psi \psi']\psi'-(\xi \psi'^{2}+\frac{\dot{V}}{g})\psi<0~~~~for~~~~r={r}_{peak}.
\end{eqnarray}

In another case that the scalar field reaches
a negative local minimum value, the relation becomes
\begin{eqnarray}\label{InfBH}
\{ \psi<0,~~~~\psi'=0~~~~and~~~~\psi''\geqslant0\}~~~~for~~~~r=r_{peak}.
\end{eqnarray}

From the relation (11) and $\xi>0$, we deduce the characteristic inequality
\begin{eqnarray}\label{BHg}
(1+\xi\psi^2)\psi''+[(1+\xi\psi^2)(\frac{2}{r}-\frac{\chi'}{2}+\frac{g'}{g})+2\xi \psi \psi']\psi'-(\xi \psi'^{2}+\frac{\dot{V}}{g})\psi>0~~~~for~~~~r=r_{peak}.
\end{eqnarray}

At the extremum point, the characteristic inequalities (10) and (12) are
in contradiction with the equation of motion (3).
It means the scalar field equation has no nonzero solution.
Then we obtain a conclusion that horizonless compact reflecting stars cannot
support static scalar fields coupled to the gradient
with non-negative coupling parameters.

\section{Conclusions}

We investigated the condensation behaviors of static
scalar fields outside asymptotically flat spherically symmetric
compact stars. At the star radius, we take the reflecting
boundary condition. In this work, we considered couplings
between scalar fields and the gradient, where this type of coupling also
appears in the next-to-leading order Ginzburg-Landau models.
In the case that coupling parameters are non-negative,
characteristic inequalities (10) and (12) are
in contradiction with the equation of motion (3),
which means that the reflecting star cannot support
static coupled scalar fields.

\begin{acknowledgments}

This work was supported by the Shandong Provincial Natural Science Foundation of China under Grant
No. ZR2018QA008. This work was supported by a grant from Qufu Normal University
of China under Grant No. xkjjc201906. This work was also supported by the Youth Innovations and Talents Project of Shandong
Provincial Colleges and Universities (Grant no. 201909118).

\end{acknowledgments}

\end{document}